# Alternating Minimization Schemes for Computing Rate-Distortion-Perception Functions with $f$-Divergence Perception Constraints

### v1

@@ Appendix B Proof of Lemma2
$\begin{split}\frac{\partial}{h(y)}\Bigg{[}-\sum_{x\in\mathcal{X}}p_{X}(x)\log\left(\sum_{y\in\mathcal{X}}h(y)A[h](x,y,s)\right)+s_{P}\sum_{y\in\mathcal{X}}p_{X}(y)\partial f\left(\frac{p_{X}(y)}{q_{Y}[h](y)}\right)+\lambda\sum_{y\in\mathcal{X}}h(y)\Bigg{]}\geq 0\end{split}$
@@ Appendix C Proof of Theorem 2
Let $\textbf{h}=(h^{(0)},h^{(1)},\ldots)$, and $\textbf{Q}=(Q_{Y|X}^{(1)},Q_{Y|X}^{(2)},\ldots)$ be the sequences of probability vectors and transition matrices obtained by the chain of alternating minimization $h^{(n)}\xrightarrow{}Q_{Y|X}^{(n+1)}\xrightarrow{}h^{(n+1)}$. Let $A^{(n)}=A[h^{(n)}]$ and define the functionals $V[h^{(n)}]$ and $W[h^{(n)}]$ as
$\begin{split}&V[h^{(n)}]=D_{KL}\left(p_{X}\cdot Q_{Y|X}^{(n+1)}||p_{X}\cdot h^{(n)}\right)+s_{D}\mathbf{E}_{Q_{Y|X}^{(n+1)}p_{X}}[d(x,y)]+s_{P}D_{f}(p_{X}||h^{(n+1)})\end{split}$ (23)
$\begin{split}&W[h^{(n)}]=s_{P}\sum_{y\in\mathcal{X}}p_{X}(y)\partial f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)-\sum_{x\in\mathcal{X}}p_{X}(x)\log\left(\sum_{y\in\mathcal{X}}h^{(n)}(y)A^{(n)}(x,,s)\right).\end{split}$ (24)
@@
$\begin{split}\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{(n)}}{Q_{Y|X}^{(n+1)}}\right)&=\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{(n)}}{h^{(n)}}\right)\\ &\quad-\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(A^{(n)}(x,,s)\right)\\ &\quad+\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\sum_{i\in\mathcal{X}}h^{(n)}(i)A^{\left(n\right)}(x,i,s)\right)\end{split}$ (25)
@@
$\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(A^{(n)}(x,,s)$ $=-s_{D}\mathbb{E}_{Q_{Y|X}^{*}}\left[d(x,y)\right]-s_{P}\sum_{y\in\mathcal{X}}h^{*}(y)f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)$
$\quad+s_{P}\sum_{y\in\mathcal{X}}h^{*}(y)\frac{p_{X}(y)}{h^{(n+1)}(y)}\partial f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right).$
@@
$\begin{split}\sum_{(x,y)\in\mathcal{X}^{2}}&p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{(n)}}{h^{(n)}}\right)-\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{*}}{h^{*}}\right)\\ &=\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{(n)}\cdot h^{*}}{h^{(n)}\cdot Q_{Y|X}^{*}}\right)\leq\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{n}\log\left(\frac{h^{*}}{h^{(n)}}\right)-1=0.\end{split}$ (26)
As a result, we obtain
$\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{(n)}}{Q_{Y|X}^{(n+1)}}\right)\leq$
$\leq\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{*}}{h^{*}}\right)+s_{D}\mathbb{E}_{Q_{Y|X}^{*}}\left[d(x,y)\right]+s_{P}D_{f}(p_{X}||h^{*})$
$\quad+s_{P}\left[\sum_{y\in\mathcal{X}}h^{*}(y)f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)-D_{f}\left(p_{X}||h^{*}\right)\right]$
$\quad-\Biggl{[}-\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\sum_{i\in\mathcal{X}}h^{(n)}(i)A^{(n)}(x,i,s)\right)+s_{P}\sum_{y\in\mathcal{X}}h^{n+1}(y)\frac{p_{X}(y)}{h^{(n+1)}(y)}\partial f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)$
$\quad+s_{P}\Biggl{[}\sum_{y\in\mathcal{X}}h^{(n+1)}(y)\frac{p_{X}(y)}{h^{(n+1)}(y)}\partial f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)-\sum_{y\in\mathcal{X}}h^{*}(y)\frac{p_{X}(y)}{h^{(n+1)}(y)}\partial f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)\Biggr{]}$
$\leq V[h^{*}]-W[h^{(n)}]+s_{P}\left[\sum_{y\in\mathcal{X}}h^{*}(y)f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)-D_{f}(p_{X}||h^{*})\right]$
$\quad+s_{P}\Biggl{[}\sum_{y\in\mathcal{X}}h^{(n+1)}(y)\frac{p_{X}(y)}{h^{(n+1)}(y)}\partial f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)-\sum_{y\in\mathcal{X}}h^{*}(y)\frac{p_{X}(y)}{h^{(n+1)}(y)}\partial f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)\Biggr{]}.$
Notice that for any iteration $n\geq 0$, we have $W[h^{(n)}]\geq V[h^{(n+1)}]\geq V[h^{*}]$ and, subsequently, the following inequalities
$0\leq W[h^{(n)}]-V[h^{*}]\leq G[h^{(n)}]$ (27)
where
$G[h^{(n)}]$ $=\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{(n+1)}}{Q_{Y|X}^{(n)}}\right)+s_{P}\left\{\mathbb{E}_{h^{*}}\left[f\left(\frac{p_{X}}{h^{(n+1)}}\right)\right]-D_{f}\left(p_{X}||h^{*}\right)\right\}$
$\quad+s_{P}\Biggl{\{}\mathbb{E}_{h^{(n+1)}}\left[\frac{p_{X}}{h^{(n+1)}}\partial f\left(\frac{p_{X}}{h^{(n+1)}}\right)\right]-\mathbb{E}_{h^{*}}\left[\frac{p_{X}}{h^{(n+1)}}\partial f\left(\frac{p_{X}}{h^{(n+1)}}\right)\right]\Biggr{\}}.$
Since $s_{P}\geq 0$ and due to the fact that
$\mathbb{E}_{h^{*}}\left[f\left(\frac{p_{X}}{h^{(n+1)}}\right)\right]-D_{f}\left(p_{X}||h^{*}\right)+\mathbb{E}_{h^{(n+1)}}\left[\frac{p_{X}}{h^{(n+1)}}\partial f\left(\frac{p_{X}}{h^{(n+1)}}\right)\right]-\mathbb{E}_{h^{*}}\left[\frac{p_{X}}{h^{(n+1)}}\partial f\left(\frac{p_{X}}{h^{(n+1)}}\right)\right]$
$=D_{f}(p_{X}||h^{(n+1)})-D_{f}(p_{X}||h^{(*)})-\sum_{i\in\mathcal{X}}\left(h^{*}(i)-h^{(n+1)}(i)\right)f\left(\frac{p_{X}(i)}{h^{(n+1)}}\right)$
$\quad+\sum_{i\in\mathcal{X}}\left(h^{*}(i)-h^{(n+1)}(i)\right)\frac{p_{X}(i)}{h^{(n+1)}}\partial f\left(\frac{p_{X}(i)}{h^{(n+1)}}\right)$
$=D_{f}\left(p_{X}||h^{(n+1)}\right)-D_{f}\left(p_{X}||h^{(*)}\right)-\sum_{i\in\mathcal{X}}\left(h^{(n+1}(i)-h^{*}(i)\right)\frac{\partial D_{f}(p_{X}||u)}{\partial u(i)}\Bigg{|}_{h^{(n+1)}}$
$=-\left(D_{f}\left(p_{X}||h^{*}\right)-T_{D_{f}(p||\cdot),h^{(n+1)}}(h^{*})\right)\stackrel{{\scriptstyle(a)}}{{\leq}}0$
where $T_{D_{f}(p||\cdot),h^{(n+1)}}(h^{*}))$ is the first order Taylor expansion of $D_{f}(p_{X}||\cdot)$, centered in $h^{(n+1)}$ and evaluated in $h^{(*)}$, and $(a)$ is verified since $D_{f}(p_{X}||\cdot)$ is a convex function in its second argument. Therefore $\forall n\in\mathbf{N}$, we obtain
$W[h^{(n)}]-V[h^{*}]\leq$ $\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{(n+1)}}{Q_{Y|X}^{(n)}}\right).$
Summing over $N$ terms, we obtain
$\sum_{n=1}^{N}(W[h^{(n)}]-V[h^{*}])$ $\leq\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\sum_{n=1}^{N}\log\left(\frac{Q_{Y|X}^{(n+1)}}{Q_{Y|X}^{(n)}}\right)$
$=\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{(N+1)}}{Q_{Y|X}^{(1)}}\right)$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{*}}{Q_{Y|X}^{(1)}}\right)$
where $(b)$ follows using the logarithm inequality Since at any iteration $n$, $W[h^{(n)}]-V[h^{*}]\geq 0$ and for all integers $N>0$ the partial sum is upper-bounded by a constant $L(h^{*},h^{(0)})$ dependent only on the initial probability assignment $h^{(0)}$, we have that $\lim_{N\to\infty}\sum_{n=1}^{N}(W[h^{(n)}]-V[h^{*}])$ exists and it is finite hence $\lim_{n\to\infty}W[h^{(n)}]-V[h^{*}]=0$. This completes the proof.
@@ Appendix D Proof of Lemma 4
By defining matrices $M$, $\Gamma$ and $C$, respectively, as in (LABEL:matrix:_Newton_M), (14), (15), the matrix $J_{T}(\cdot)$ can be rewritten as in (13). To prove its invertibility, we need to ensure that 00 is not part of the set of eigenvalues of $J_{T}(y)$, i.e., $0\notin\operatorname{eig}(J_{T}(y)),~{}\forall y\in\mathbb{R}^{|\mathcal{X}|}$. Noticing that $M_{i,j}\geq 0,\forall(i,j)\in\mathcal{X}^{2}$ and $\sum_{i\in\mathcal{X}}M_{i,i}(y)=C_{i}(y)$, we can define the sets $D_{i}$ as:
$D_{i}=\{\lambda\in\mathbb{R}:\left|\lambda-\left(C[h^{(n)},u](i)-M[h^{(n)},u](i,i)\right)\right|\leq C[h^{(n)},u](i)-M[h^{(n)},u](i,i)\}$ (28)
@@
$J_{T}[h^{(n)},u]=I+\left(C[h^{(n)},u]-M[h^{(n)},u]\right)\cdot\Gamma[h^{(n)},u]\geq I>0$
@@ Appendix E Proof of Theorem 4
Let $V[\cdot]$, $W[\cdot]$ be the functionals defined in (23) and (24), respectively. Moreover, let $\hat{W}[\cdot]$ be a functional obtained by the alternating sequence $\hat{h}^{(n)}\to\hat{Q}_{Y|X}^{(n+1)}\to\hat{h}^{(n+1)}$ substituting $\hat{Q}_{Y|X}$ with fixed $\hat{h}$ as follows
$\begin{split}\hat{W}[\hat{h}^{(n)}]=&-\sum_{x\in\mathcal{X}}p_{X}(x)\log\left(\sum_{y\in\mathcal{X}}\hat{h}^{(n)}(y)A^{(n)}(x,,s)\right)+s_{P}\sum_{y\in\mathcal{X}}p_{X}(y)\frac{p_{X}(y)}{v^{(n)}(y)}\partial f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\\ &+s_{P}\Bigg{[}\sum_{y\in\mathcal{X}}\hat{h}^{(n+1)}f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)-\sum_{y\in\mathcal{X}}\hat{h}^{(n+1)}f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\Bigg{]}\end{split}$ (29)
where $A^{(n)}=A[\hat{h}^{(n)}]$. Similarly to Theorem 2, we let $h^{*}$ be any probability vector and $Q_{Y|X}^{*}=Q_{Y|X}[h^{*}]$ such that $R(D,P)=V[h^{*}]-s_{D}D-s_{P}P$ and consider that
$\begin{split}\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{\hat{Q}_{Y|X}^{(n)}}{\hat{Q}_{Y|X}^{(n+1)}}\right)&=\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{\hat{Q}_{Y|X}^{(n)}}{\hat{h}^{(n)}}\right)-\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(A^{(n)}(x,,s)\right)\\ &\quad+\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\sum_{i\in\mathcal{X}}\hat{h}^{(n)}(i)A^{(n)}(x,i,s)\right).\end{split}$
@@
$\begin{split}&\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{\hat{Q}_{Y|X}^{(n)}}{\hat{Q}_{Y|X}^{(n+1)}}\right)\leq\quad V[h^{*}]-\hat{W}[\hat{h}^{(n)}]\\ &\quad+s_{P}\Bigg{[}\sum_{y\in\mathcal{X}}h^{*}(y)f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)-D_{f}(p_{X}||h^{*})+\sum_{y\in\mathcal{X}}h^{(n+1)}(y)\frac{p_{X}(y)}{v^{(n)}(y)}\partial f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\\ &\quad\qquad-\sum_{y\in\mathcal{X}}h^{*}(y)\frac{p_{X}(y)}{v^{(n)}(y)}\partial f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)+\sum_{y\in\mathcal{X}}\hat{h}^{(n+1)}f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)-\sum_{y\in\mathcal{X}}\hat{h}^{(n+1)}f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\Bigg{]}.\end{split}$ (30)
@@
$\begin{split}&\sum_{y\in\mathcal{X}}h^{*}(y)f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)-D_{f}\left(p_{X}||h^{*}\right)+\sum_{y\in\mathcal{X}}h^{(n+1)}(y)\frac{p_{X}(y)}{v^{(n)}(y)}\partial f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\\ &\quad-\sum_{y\in\mathcal{X}}h^{*}(y)\frac{p_{X}(y)}{v^{(n)}(y)}\partial f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)+\sum_{y\in\mathcal{X}}\hat{h}^{(n+1)}f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)-\sum_{y\in\mathcal{X}}\hat{h}^{(n+1)}f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\\ &=D_{f}\left(p_{X}||v^{(n)}\right)-D_{f}\left(p_{X}||h^{*}\right)+\sum_{i\in\mathcal{X}}(h^{*}-v^{(n)})\left[f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)-\frac{p_{X}(y)}{v^{(n)}(y)}\partial f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\right]\\ &\quad+D_{f}\left(p_{X}||h^{(n+1)}\right)-D_{f}\left(p_{X}||v^{(n)}\right)+\sum_{i\in\mathcal{X}}\left(v^{(n)}-h^{(n+1)}\right)\left[f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)-\frac{p_{X}(y)}{v^{(n)}(y)}\partial f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\right]\\ &=-\left[D_{f}\left(p_{X}||h^{*}\right)-T_{D_{f}(p||\cdot),v^{(n)}}(h^{*})\right]+\left[D_{f}(p_{X}||h^{(n+1)})-T_{D_{f}(p||\cdot),v^{(n)}}\left(h^{(n+1)}\right)\right]\\ &\stackrel{{\scriptstyle(a)}}{{\leq}}D_{f}\left(p_{X}||h^{(n+1)}\right)-T_{D_{f}(p||\cdot),v^{(n)}}\left(h^{(n+1)}\right)\end{split}$
@@
$\begin{split}G[\hat{h}^{(n)}]=\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{\hat{Q}_{Y|X}^{(n+1)}}{\hat{Q}_{Y|X}^{(n)}}\right)+s_{P}\left[D_{f}\left(p_{X}||h^{(n+1)}\right)-T_{D_{f}(p||\cdot),v^{(n)}}\left(h^{(n+1)}\right)\right].\end{split}$
@@
$\begin{split}\sum_{n=1}^{N}\hat{W}[\hat{h}^{(n)}]-V[h^{*}]\leq&\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\sum_{n=1}^{N}\log\left(\frac{\hat{Q}_{Y|X}^{(n+1)}}{\hat{Q}_{Y|X}^{(n)}}\right)\\ &+s_{P}\sum_{n=1}^{N}D_{f}\left(p_{X}||h^{(n+1)}\right)-T_{D_{f}(p||\cdot),v^{(n)}}\left(h^{(n+1)}\right)\\ \leq&\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{*}}{\hat{Q}_{Y|X}^{(1)}}\right)+\sum_{n=1}^{N}o\left(||h^{(n+1)}-v^{(n)}||\right)\\ \leq&\quad\tilde{L}(h^{*},h^{(0)})\end{split}$
where $\tilde{L}(h^{*},h^{(0)})$ is finite if the limit $\lim_{n\to\infty}||h^{(n+1)}-v^{(n)}||=0$ converges at least linearly. Thus, we can rewrite

### v2

@@ Appendix B Proof of Lemma2
$\begin{split}\frac{\partial}{h(y)}&\Bigg{[}-\sum_{x\in\mathcal{X}}p_{X}(x)\log\left(\sum_{y\in\mathcal{X}}h(y)A[h](x,y,s)\right)\\ &+s_{P}\sum_{y\in\mathcal{X}}p_{X}(y)\partial f\left(\frac{p_{X}(y)}{q_{Y}[h](y)}\right)+\lambda\sum_{y\in\mathcal{X}}h(y)\Bigg{]}\geq 0\end{split}$
@@ Appendix C Proof of Theorem 2
Let $\textbf{h}=(h^{(0)},h^{(1)},\ldots)$, and $\textbf{Q}=(Q_{Y|X}^{(1)},Q_{Y|X}^{(2)},\ldots)$ be the sequences of probability vectors and transition matrices obtained by the chain of alternating minimization $h^{(n)}\xrightarrow{}Q_{Y|X}^{(n+1)}\xrightarrow{}h^{(n+1)}$. Let $A^{(n)}=A[h^{(n+1)}]$ and define the functionals $V[h^{(n)}]$ and $W[h^{(n)}]$ as
$V[h^{(n)}]=$ $D_{KL}\left(p_{X}\cdot Q_{Y|X}^{(n+1)}||p_{X}\cdot h^{(n)}\right)$ (24)
$+s_{D}\mathbb{E}_{p_{X}Q_{Y|X}^{(n+1)}}\left[d(X,Y)\right]+s_{P}D_{f}(p_{X}||h^{(n+1)})$
$W[h^{(n)}]$ $=s_{P}\sum_{y\in\mathcal{X}}p_{X}(y)\partial f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)$ (25)
$~~~~-\sum_{x\in\mathcal{X}}p_{X}(x)\log\left(\sum_{y\in\mathcal{X}}h^{(n)}(y)A^{(n)}(x,,s)\right)$
@@
$\begin{split}&\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{(n)}}{Q_{Y|X}^{(n+1)}}\right)\\ &=\quad\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{(n)}}{h^{(n)}}\right)\\ &\qquad-\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(A^{(n)}(x,,s)\right)\\ &\qquad+\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\sum_{i\in\mathcal{X}}h^{(n)}(i)A^{\left(n\right)}(x,i,s)\right)\end{split}$ (26)
@@
$\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(A^{(n)}(x,,s)$
$\quad=-s_{D}\mathbb{E}_{Q_{Y|X}^{*}}\left[d(X,Y)\right]-s_{P}\sum_{y\in\mathcal{X}}h^{*}(y)f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)$
$\qquad+s_{P}\sum_{y\in\mathcal{X}}h^{*}(y)\frac{p_{X}(y)}{h^{(n+1)}(y)}\partial f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right).$
@@
$\begin{split}&\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{(n)}}{h^{(n)}}\right)-p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{*}}{h^{*}}\right)\\ &\quad=\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{Q_{Y|X}^{(n)}\cdot h^{*}}{h^{(n)}\cdot Q_{Y|X}^{*}}\right)\\ &\quad\leq\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{(n)}\log\left(\frac{h^{*}}{h^{(n)}}\right)-1=0.\end{split}$ (27)
As a result, we obtain Since at any iteration $n$, $W[h^{(n)}]-V[h^{*}]\geq 0$ and for all integers $N>0$ the partial sum is upper-bounded by a constant $L(h^{*},h^{(0)})$ dependent only on the initial probability assignment $h^{(0)}$, we have that $\lim_{N\to\infty}\sum_{n=1}^{N}(W[h^{(n)}]-V[h^{*}])$ exists and it is finite hence $\lim_{n\to\infty}W[h^{(n)}]-V[h^{*}]=0$. This completes the proof.
@@ Appendix D Proof of Lemma 4
By defining matrices $M$, $\Gamma$ and $C$, respectively, as in (14), (15), (16), the matrix $J_{T}(\cdot)$ can be rewritten as in (13). To prove its invertibility, we need to ensure that 0 is not part of the set of eigenvalues of $J_{T}(y)$, i.e., $0\notin\operatorname{eig}(J_{T}(y)),~\forall y\in\mathbb{R}^{|\mathcal{X}|}$. Noticing that $M[h^{(n)},u](i,j)\geq 0,\forall(i,j)\in\mathcal{X}^{2}$ and $\sum_{i\in\mathcal{X}}M[h^{(n)},u](i,j)=C[h^{(n)},u](i)$, we can define the sets $D_{i}$ as:
$D_{i}=\Big{\{}\lambda\in\mathbb{R}:$ $\left|\lambda-\left(C[h^{(n)},u](i)-M[h^{(n)},u](i,i)\right)\right|$
$\leq C[h^{(n)},u](i)-M[h^{(n)},u](i,i)\Big{\}}$
@@
$J_{T}[h^{(n)},u]$ $=I+\left(C[h^{(n)},u]-M[h^{(n)},u]\right)\cdot\Gamma[h^{(n)},u]$
$\geq I>0$
@@ Appendix E Proof of Theorem 4
Let $V[\cdot]$, $W[\cdot]$ be the functionals defined in (24) and (25), respectively. Moreover, let $\hat{W}$$\cdot$] be a functional obtained by the alternating sequence $\hat{h}^{(n)}\to\hat{Q}_{Y|X}^{(n+1)}\to\hat{h}^{(n+1)}$ substituting $\hat{Q}_{Y|X}$ with fixed $\hat{h}$ as follows
where $A^{(n)}=A[v^{(n)}]$. Similarly to Theorem 2, we let $h^{*}$ be any probability vector and $Q_{Y|X}^{*}=Q_{Y|X}[h^{*}]$ such that $R(D,P)=V[h^{*}]-s_{D}D-s_{P}P$ and consider that
$\begin{split}&\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{\hat{Q}_{Y|X}^{(n)}}{\hat{Q}_{Y|X}^{(n+1)}}\right)\\ &\quad=\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{\hat{Q}_{Y|X}^{(n)}}{\hat{h}^{(n)}}\right)\\ &\qquad-\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(A^{(n)}(x,,s)\right)\\ &\qquad+\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\sum_{i\in\mathcal{X}}\hat{h}^{(n)}(i)A^{(n)}(x,i,s)\right).\end{split}$
@@
$\begin{split}&\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{\hat{Q}_{Y|X}^{(n)}}{\hat{Q}_{Y|X}^{(n+1)}}\right)\leq\quad V[h^{*}]-\hat{W}[\hat{h}^{(n)}]\\ &\quad+s_{P}\Bigg{[}\sum_{y\in\mathcal{X}}h^{*}(y)f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)-D_{f}(p_{X}||h^{*})\\ &\qquad~~~~~~+\sum_{y\in\mathcal{X}}h^{(n+1)}(y)\frac{p_{X}(y)}{v^{(n)}(y)}\partial f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\\ &\qquad~~~~~~-\sum_{y\in\mathcal{X}}h^{*}(y)\frac{p_{X}(y)}{v^{(n)}(y)}\partial f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\\ &\qquad~~~~~~+\sum_{y\in\mathcal{X}}\hat{h}^{(n+1)}f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)\\ &\qquad~~~~~~-\sum_{y\in\mathcal{X}}\hat{h}^{(n+1)}f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\Bigg{]}.\end{split}$ (29)
@@
$\begin{split}&\sum_{y\in\mathcal{X}}h^{*}(y)f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)-D_{f}\left(p_{X}||h^{*}\right)\\ &\quad+\sum_{y\in\mathcal{X}}h^{(n+1)}(y)\frac{p_{X}(y)}{v^{(n)}(y)}\partial f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\\ &\quad-\sum_{y\in\mathcal{X}}h^{*}(y)\frac{p_{X}(y)}{v^{(n)}(y)}\partial f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\\ &\quad+\sum_{y\in\mathcal{X}}\hat{h}^{(n+1)}f\left(\frac{p_{X}(y)}{h^{(n+1)}(y)}\right)-\sum_{y\in\mathcal{X}}\hat{h}^{(n+1)}f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\\ &=D_{f}\left(p_{X}||v^{(n)}\right)-D_{f}\left(p_{X}||h^{*}\right)\\ &\quad+\sum_{i\in\mathcal{X}}(h^{*}-v^{(n)})\left[f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)-\frac{p_{X}(y)}{v^{(n)}(y)}\partial f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\right]\\ &\quad+D_{f}\left(p_{X}||h^{(n+1)}\right)-D_{f}\left(p_{X}||v^{(n)}\right)\\ &\quad+\sum_{i\in\mathcal{X}}\left(v^{(n)}-h^{(n+1)}\right)f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\\ &\quad-\sum_{i\in\mathcal{X}}\left(v^{(n)}-h^{(n+1)}\right)\frac{p_{X}(y)}{v^{(n)}(y)}\partial f\left(\frac{p_{X}(y)}{v^{(n)}(y)}\right)\\ &=-\left[D_{f}\left(p_{X}||h^{*}\right)-T_{D_{f}(p||\cdot),v^{(n)}}(h^{*})\right]\\ &\quad+\left[D_{f}(p_{X}||h^{(n+1)})-T_{D_{f}(p||\cdot),v^{(n)}}\left(h^{(n+1)}\right)\right]\\ &\stackrel{{\scriptstyle(a)}}{{\leq}}D_{f}\left(p_{X}||h^{(n+1)}\right)-T_{D_{f}(p||\cdot),v^{(n)}}\left(h^{(n+1)}\right)\end{split}$
@@
$\begin{split}&G[\hat{h}^{(n)}]=\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}^{*}\log\left(\frac{\hat{Q}_{Y|X}^{(n+1)}}{\hat{Q}_{Y|X}^{(n)}}\right)\\ &\qquad+s_{P}\left[D_{f}\left(p_{X}||h^{(n+1)}\right)-T_{D_{f}(p||\cdot),v^{(n)}}\left(h^{(n+1)}\right)\right].\end{split}$
@@
[…]